The 7 percent bonds issued by Modern Kitchens pay interest semiannually, mature in eight years, and have a $1,000 face value. Currently, the yield to maturity for these bonds is 7.22%. What is the market price per bond?

$986.81 Please show work
Answer #1
Price of bond is the present value of cash flows from bond.
Semi annual coupon interest = Face Value * Semi annual coupon rate
= $        1,000 * 3.50%
= $              35
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.0361)^-16)/0.0361 i 7.22%/2 = 0.0361
= 11.994816 n 8*2 = 16
Present value of 1 = (1+i)^-n
= (1+0.0361)^-16
= 0.5669872
Present Value of coupon interest = Semi annual coupon interest*Present value of annuity of 1
= $              35 * 11.99482
= $     419.82
Present value of face Value = Face Value * Present value of 1
= $        1,000 * 0.566987
= $     566.99
Present Value of Cash flows = Present Value of coupon interest + Present value of face Value
= $     419.82 + $ 566.99
= $     986.81
Present value of cash flows from bond is $ 986.81.So, Price of Bond is $ 986.81